OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 3y = 3 3x – 2y = 11 What is the solution to the system? (–3, 3) (–1, –7) (1, –4) (3, –1)

OpenStudy (phi):

Each equation is the equation of a line. They want you to plot the lines (find 2 points on a line and connect the dots) The answer is the spot where the lines cross

OpenStudy (nubeer):

(3,-1)

OpenStudy (anonymous):

i tried all of them only one i could get was 3,-1 but only the 3 worked and not the -1 :(

OpenStudy (anonymous):

on the second one

OpenStudy (phi):

i tried all of them only one i could get was 3,-1 but only the 3 worked and not the -1 ?? if you replace x with 3 and y with -1 in each equation: 2x + 3y = 3 2*3 + 3*-1= 3 it should work same for the other equation

OpenStudy (phi):

now try the 2nd equation 3x – 2y = 11 with x= 3 and y= -1